OpenStudy (anonymous):

Handy Mandy fixes major appliances such as refrigirator and air conditioners . She charges a fixed fee for a home visit plus an hourly rate for the repair work. Use the info. Below Hours : 3. 7. 10 Total cost: $97. $205. $286 A) what is Mandy's fixed fee? B) what is Mandy's hourly rate ?

OpenStudy (texaschic101):

x = fixed fee y = hourly rate x + 3y = 97 --- multiply by -1 x + 7y = 205 --------------- -x - 3y = -97 (result of multiplying by -1) x + 7y = 205 ---------------add 4y = 108 y = 27 x + 3y = 97 x + 3(27) = 97 x + 81 = 97 x = 97 - 81 x = 16 check... x + 7y = 205 16 + 7(27) = 205 16 + 189 = 205 205 = 205 (correct) check.. x + 10y = 286 16 + 10(27) = 286 16 + 270 = 286 286 = 286 (correct) so the standard fee (x) = 16 and the hourly rate (y) = 27

OpenStudy (texaschic101):

oops...I meant fixed fee, not standard fee

OpenStudy (anonymous):

Thanksssssss